# Rivers under Noise

### v1

@@ Abstract
We consider the deterministic and stochastic versions of a first order non-autonomous differential equation which allows us to discuss the persistence of rivers (“fleuves”) under noise.
Keywords: SDEs, rivers, trichotomy
AMS2020 classification: 34F05, 37H05, 37H30, 60H10.
@@ 1 Introduction
This paper uses a simple example in the pursuit of understanding the stochastic counterpart of the river (“fleuve”) phenomenon.
@@
The present paper is devoted to an analysis of a simple example of the stochastic counterpart of a polynomial non-autonomous ordinary differential equation which exhibits rivers in which a trichotomy of asymptotic fates of solutions occurs. The structure of the paper is as follows. In Section 2 we discuss the structure of solutions of the deterministic version of our equation and exhibit the asymptotics of the critical solution; this material is standard and helped by the existence of an explicit form of the solutions. In Section 3, we briefly study the much simpler (deterministic and random) unstable linear case in which there is a repelling river of sub-exponential growth. It is easy to analyze since there is an explicit formula for the solution and for the repelling river. In Section 4, we state our main result: the non-autonomous quadratic stochastic differential equation (4.1) also admits a trichotomy: there is a random repelling river with the property that trajectories starting above it blow up in finite time while trajectories starting below the repelling river converge to 0 and the river which separates the two regimes has asymptotic linear growth (like in the deterministic case). Note that, contrary to the linear case, there is an attracting solution (any solution which converges to 0 attracts all trajectories starting below the repelling river). We formulate our results as four theorems which we prove in the subsequent sections. Our main tools are well-known estimates of exit probabilities of diffusions from intervals of the real line which we quote in the appendix for the reader’s convenience. Finally, in Section 8, we discuss extending other tools available in the deterministic setting (Ważewski principle, asymptotic expansions) to the stochastic context.
@@ 2 The deterministic situation
In the deterministic case, the equation
(2.1) $x^{\prime}=x^{2}-tx,\quad t>0,\quad x(0)=x_{0},$
is easy to analyse as the solution can be found explicitly,
(2.2) $x(t)=\frac{2x_{0}\exp\left(-\frac{t^{2}}{2}\right)}{2-x_{0}\sqrt{2\pi}\,\hbox{erf}\,\left(\frac{t}{\sqrt{2}}\right)}.$
Here erf is the error function,
@@
$\lim_{z\rightarrow\infty}\hbox{erf}\,(z)=1$, from which we immediately obtain the following trichotomy in positive time : If $x_{0}>\sqrt{2/\pi}$, solutions blow up in finite time; if $x_{0}=\sqrt{2/\pi}$, $x(t)\rightarrow\infty$ as $t\rightarrow\infty$, and finally if $x_{0}<\sqrt{2/\pi}$, $x(t)\rightarrow 0$ as $t\rightarrow\infty$. We call the solution (2.2) with $x_{0}=\sqrt{2/\pi}$ the critical solution and denote it by $x_{c}(t)$.
@@
Given the explicit form, the asymptotics of the critical solution as $t\rightarrow\infty$ is easily computed; it is given by
(2.3) $x_{c}(t)\sim t+\frac{1}{t}-\frac{2}{t^{3}}+O\left(\frac{1}{t^{5}}\right)$
Note that in terms of rivers, $x_{c}(t)$ provides a repelling river, while $x(t)\equiv 0$ and all the positive semi-orbits that converge to it, constitute the attracting river.
The asymptotic expansion in (2.3) can be obtained directly by finding the asymptotic expansion of the repelling river as explained in [1].
@@ 3 The linear case
$\text{\rm{d}}X(t)=\big{(}cX(t)+f(t)\big{)}\,\text{\rm{d}}t+\sigma\text{\rm{d}}W(t),\;t\geq 0,$
@@
The SDE with initial condition $X(0)=x$ has the explicit solution
@@
It follows that all trajectories are exponentially unstable with $\frac{\text{\rm{d}}}{\text{\rm{d}}x}X(t,x)=\text{\rm{e}}^{ct},\,t\geq 0,\,x\in\mathbb{R}$. We now assume that
@@
Denote the unique solution $X(t),\,t\geq 0$ with initial condition
@@
by $\mathscr{R}(t),\,t\geq 0$ ($\mathscr{R}$ stands for river). Then
$\mathscr{R}(t)=-\text{\rm{e}}^{ct}\int_{t}^{\infty}\text{\rm{e}}^{-cs}f(s)\,\text{\rm{d}}s-\sigma\text{\rm{e}}^{ct}\int_{t}^{\infty}\text{\rm{e}}^{-cs}\,\text{\rm{d}}W(s)=-\text{\rm{e}}^{ct}\int_{t}^{\infty}\text{\rm{e}}^{-cs}f(s)\,\text{\rm{d}}s-\sigma A(t),$
@@
so solutions grow (respectively decrease) exponentially with rate $c$ when $x$ is above (respectively below) $\mathscr{R}(0)$ while $\big{|}\mathscr{R}(t)\big{|}=o\big{(}\text{\rm{e}}^{ct}\big{)}$ when $x=\mathscr{R}(0)$. In particular, there is a trichotomy concerning the asymptotic behaviour just like in the deterministic case ($\sigma\equiv 0$) with $f\equiv 0$ in which the repelling river is $X(t,x)\equiv 0$.
@@ 4 Main results
We now study the additive noise stochastic equivalent of (2.1) namely the stochastic differential equation
(4.1) $\text{\rm{d}}X(t)=X(t)\big{(}X(t)-t\big{)}\,\text{\rm{d}}t+\sigma\,\text{\rm{d}}W(t),\,X(s)=x,\,t\geq s,$
@@
Before stating our first main result, we collect some elementary properties of the following more general model:
(4.2) $X(t)=x+\int_{s}^{t}X(u)\big{(}X(u)-u\big{)}\,\text{\rm{d}}u+H(t)-H(s),\,t\geq s,$
where $H:[0,\infty)\to\mathbb{R}$ is a given (random or deterministic) continuous function satisfying $H(0)=0$, $x\in\mathbb{R}$, and $s\geq 0$. Again we denote the unique local solution of (4.2) by $X_{s,t}(x)$ and define the blow-up time by
@@
Note that for each $x\in\mathbb{R}$ and $v\geq s\geq 0$, we have
$\inf_{s\leq t\leq v}X_{s,t}(x)\geq x\wedge 0+\inf_{s\leq t\leq u\leq v}\big{(}H(u)-H(t)\big{)}$
since $y(y-r)\geq 0$ for $r\geq 0$ and $y\leq 0$. In particular, the function $t\mapsto X_{s,t}(x)$ cannot reach $\infty$- ∞ in finite time. Further, if $X_{s,t}(x)\geq A$ for some $t\in[s,v]$ and $A\geq v+\inf_{s\leq\kappa\leq u\leq v}\big{(}H(u)-H(\kappa)\big{)}$, then
@@
since $y(y-r)\geq 0$ for $y\geq v$ and $r\in[s,v]$. This implies that $\lim_{t\uparrow\beta_{s}(x)}X_{s,t}(x)=\infty$ whenever $\beta_{s}(x)<\infty$. The function $x\mapsto\beta_{s}(x)$ is non-decreasing for each $s\geq 0$.
@@
The following monotonicity property will be important. Its proof is straightforward.
Lemma 4.1.
$x<y$ implies $X_{s,t}(x)\leq X_{s,t}(y)$ for $t\geq s$ and even $X_{s,t}(x)<X_{s,t}(y)$ when $\beta_{s}(x)>t$.
For any $s\geq 0$, we have $\beta_{s}(x)<\infty$ for sufficiently large $x$. To see this, let $H_{*}$ be a lower bound of $H(t)-H(s)$ for $t\in[s,s+1]$. Then $\beta_{s}(x)<s+1$ for sufficiently large $x$. By monotonicity, the set $I_{1}(s)$ of all $x\in\mathbb{R}$ for which $\beta_{s}(x)<\infty$ is an interval which is unbounded above and non-empty. The function
$\mathscr{R}(s):=\inf\{x\in\mathbb{R}:\,\beta_{s}(x)<\infty\}\in[-\infty,\infty)$
will play an important role. By continuity of solutions, $\mathscr{R}(t)$ is a particular solution of equation (4.2) on the set where $\mathscr{R}$ is larger than $\infty$- ∞. $\mathscr{R}$ does not blow up in finite time: assume it blows up at time $T<\infty$, then a solution starting at $(T,x)$ will blow up when $x$ is sufficiently large contradicting the definition of $\beta_{T}$. In particular, the set $I_{1}(s)$ is open (depending on $H$ it may be all of $\mathbb{R}$ or a proper (non-empty) subset of $\mathbb{R}$).
@@
Theorem 4.2 implies that for each $s\geq 0$ and $x\in\mathbb{R}$ the set $\Omega_{s,x}\subset\Omega$ consisting of those $\omega\in\Omega$ for which the solution starting at $(s,x)$ neither blows up nor converges to zero has measure 0. For $s\geq 0$ denote the union of all $\Omega_{s,x},\,x\in\mathbb{R}$ by $\Omega_{s}$. Since this is an uncountable union it does not follow that $\Omega_{s}$ has measure 0. Indeed we will see in the following that $\mathbb{P}(\Omega_{s})>0$ for every $s\geq 0$. In order to obtain a better understanding of the set $\Omega_{s}$ consider the set $I_{1}(s,\omega)$ of those $x\in\mathbb{R}$ for which $\beta_{s}(x)<\infty$ (as above). We saw that $I_{1}(s,\omega)$ is a non-empty open interval which is unbounded from above. Denote the set of $x$ for which $X_{s,t}(x)$ converges to 0 by $I_{2}(s,\omega)$. Again, by monotonicity of solutions, $I_{2}(s,\omega)$ is a (possibly empty) interval. Theorem 4.2 implies that every $x\in\mathbb{R}$ is almost surely contained in the (disjoint) union of these two intervals, so the complement of the union of these intervals can contain at most one point. Recall that $\mathscr{R}(s,\omega)$ is the infimum of $I_{1}(s,\omega)$ (which necessarily coincides with the supremum of $I_{2}(s,\omega)$). Note that $\mathscr{R}(s,\omega)=-\infty$ iff $I_{2}(s,\omega)$ is empty iff $I_{1}(s,\omega)=\mathbb{R}$. We know that $\mathscr{R}(s,\omega)$ does not belong to $I_{1}(s,\omega)$ but at the moment, it is unclear if $\mathscr{R}(s,\omega)$ belongs to $I_{2}(s,\omega)$. We will see later that it almost surely doesn’t. Therefore (almost surely), $\omega\in\Omega_{s}$ iff $\mathscr{R}(s,\omega)>-\infty$ and $\omega\in\Omega_{s,x}$ iff $\mathscr{R}(s,\omega)=x$. In particular, $\mathbb{P}(\Omega_{s})=\mathbb{P}(\mathscr{R}(s,\omega)>-\infty)$ and we will see in Theorem 4.6 that this quantity converges to 1 as $s\to\infty$.
@@
Recall that $\mathscr{R}(s)\in[-\infty,\infty)$ for all $s\geq 0$ and that $\mathscr{R}(s)=-\infty$ implies $\mathscr{R}(t)=-\infty$ for every $t<s$. Further, $X_{s,t}(\mathscr{R}(s))=\mathscr{R}(t)$ whenever $t\geq s$ and $\mathscr{R}(s)>-\infty$, so $\mathscr{R}(t)$ is a solution of (4.1) on the interval where it is larger than $\infty$- ∞. Theorem 4.4 implies that, for each $\varepsilon>0$ and $\alpha\in(0,1/2)$, $\lim_{s\to\infty}\mathbb{P}\big{(}s^{\alpha}|\mathscr{R}(s)-s|\geq\varepsilon\big{)}=0$ but the following result is stronger.
@@
In the deterministic case, it follows from the fact that the initial value of the repelling river at time 0 is positive, together with the fact that the derivative is zero on the diagonal $x=t$ that the function $\mathscr{R}$ lies above the diagonal, i.e. $\mathscr{R}(s)>s$ for all $s\geq 0$. This is not true in the stochastic case.
@@
(4.3) $\big{|}\mathscr{R}_{n}(t)-\mathscr{R}(t)\big{|}=o(t^{-\frac{3}{2}n-\frac{1}{2}+\varepsilon}),\;t\to\infty,$
@@
Define $\Gamma_{n}(t):=\sup_{u\geq t}\big{|}\mathscr{R}_{n}(u)-\mathscr{R}(u)\big{|}$. By Theorem 4.6, there exists some $t_{0}(\varepsilon,\omega)<\infty$ such that $|Z(u)|\leq u^{-\frac{1}{2}+\varepsilon}$ for all $u\geq t_{0}(\varepsilon,\omega)$. Hence, for $t\geq t_{0}(\varepsilon,\omega)$,
@@
Let $t\geq t_{1}(\varepsilon,\omega):=3\vee t_{0}(\varepsilon,\omega)$ and assume without loss of generality that $\varepsilon<1/2$. Then, by induction, $\Gamma_{n}(t)\leq t^{-\frac{1}{2}+\varepsilon}$ (this is true for $n=1$) since
@@
  • The solution with initial condition $(s,x)$ with $x\in(0,s-1]$ reaches the set $\{0,s\}$ in finite time. Let $\rho(s,x)$ be the probability that it hits level 0 before level $s$. Lemma 5.7 says that $\lim_{s\to\infty}\sup_{x\leq s}\rho(s,x)=1$ which implies that $\overline{\rho}(s):=\inf_{t\geq s}\inf_{x\in(0,s-1]}\rho(t,x)$ converges to 1 as well.
  • The solution with initial condition $(s,x)$, $x<0$ will hit 0 after finite time (irrespective of the starting time $s\geq 0$). This is Lemma 5.6.
@@
These statements together imply the claims in Theorems 4.2 and 4.4 as follows:
@@
(5.1) $\inf_{x\in\mathbb{R}}\Big{(}B(s,x)+C(s,x)\Big{)}\geq\overline{B}(s)\wedge\big{(}\overline{\rho}(s)\overline{\chi}(s)\big{)}$
@@
(5.2) $\kappa(s,x)\leq\int_{\mathbb{R}}\kappa(t,y)\,\mathbb{P}\big{(}X_{s,t}(x)\in\text{\rm{d}}y\big{)}$
@@
We will now formulate and prove the lemmas mentioned above. Since the proof of Theorem 4.6 will require finer estimates of $B(s,x)$ than those of Theorems 4.2 and 4.4 we will formulate the lemmas accordingly. The first lemma provides a lower bound for the blow-up probability when a trajectory starts slightly above the diagonal. The stated result is in fact more precise than needed to prove the two theorems. It will be convenient to work with the process
@@
(5.3) $\text{\rm{d}}Y(t)=\big{(}Y(t)(Y(t)+t)-1\big{)}\,\text{\rm{d}}t+\sigma\,\text{\rm{d}}W(t),\,t\geq s,\,Y(s)=x.$
@@
$B\big{(}s,s+s^{-\alpha}\big{)}\geq\exp\Big{\{}-\frac{1}{\sigma^{2}}s^{1-2\alpha}\big{(}1+o(1)\big{)}\Big{\}},\,s\to\infty.$
@@
(5.4) $\text{\rm{d}}\,\overline{Y}(t)=\big{(}\overline{Y}(t)(\overline{Y}(t)+s)-1\big{)}\,\text{\rm{d}}t+\sigma\,\text{\rm{d}}W(t),\,t\geq s,\overline{Y}(s)=\gamma.$
$\overline{Y}$ solves an equation with coefficients independent of $t$ and is therefore easier to analyze than $Y$. Note that $\overline{Y}(t)\leq Y(t)$ as long as $Y$ is non-negative. Further, $\overline{Y}$ blows up to $\infty$ with probability 1 by Feller’s test of explosion, see Lemma 9.1c), applied to $(l,r)=\mathbb{R}$. Therefore,

### v2

# Rivers under Additive Noise

@@ Abstract
We consider the deterministic and stochastic versions of a first order non-autonomous differential equation which allows us to discuss the persistence of rivers (“fleuves”) under additive noise.
keywords:
stochastic differential equations, rivers, trichotomy
pacs:
MSC Classification]34F05, 37H05, 37H30, 60H10
@@ 1 Introduction
This paper uses a simple example in the pursuit of understanding the stochastic counterpart of the river (“fleuve”) phenomenon. We believe that this example contains all the important ingredients of the general theory.
@@
The present paper is devoted to an analysis of a simple example of the stochastic counterpart of a polynomial non-autonomous ordinary differential equation which exhibits rivers in which a trichotomy of asymptotic fates of solutions occurs. The structure of the paper is as follows. In Section 2 we discuss the structure of solutions of the deterministic version of our equation (2.1), and exhibit the asymptotics of the critical solution; this material is standard and is helped by the existence of an explicit formula for the general solution. In Section 3, we briefly study the much simpler (deterministic and random) unstable linear case in which there is a repelling river of sub-exponential growth. This case is easy to analyze since again there is an explicit formula for the general solution and for the repelling river. In Section 4, we state our main result: the non-autonomous quadratic stochastic differential equation (4.1), the additive noise counterpart of (2.1) also admits a trichotomy: there is a random repelling river with the property that trajectories starting above it blow up in finite time while trajectories starting below the repelling river converge to 0 and the river which separates the two regimes has linear growth as $t\rightarrow\infty$ (as in the deterministic case). Note that, contrary to the linear case, there is an attracting solution (any solution which converges to 0 attracts all trajectories starting below the repelling river). We formulate our results as four theorems which we prove in the subsequent sections. Our main tools are well-known estimates of exit probabilities of diffusions from intervals of the real line which we quote in the appendix for the reader’s convenience. Finally, in Section 8, we discuss extending other tools available in the deterministic setting (Ważewski principle, asymptotic expansions) to the stochastic context.
@@ 2 The deterministic situation
The deterministic equation we start with is
$x^{\prime}=x^{2}-tx,\quad t>0,\quad x(0)=x_{0}.$ (2.1)
It is easy to analyse as the general solution of the initial value problem can be found explicitly,
$x(t)=\frac{2x_{0}\exp\left(-\frac{t^{2}}{2}\right)}{2-x_{0}\sqrt{2\pi}\,\hbox{erf}\,\left(\frac{t}{\sqrt{2}}\right)}.$ (2.2)
Here $\hbox{erf}(\cdot)$ is the error function,
@@
$\lim_{z\rightarrow\infty}\hbox{erf}\,(z)=1$, from which we immediately obtain the following trichotomy of fates of solutions in positive time : If $x_{0}>\sqrt{2/\pi}$, solutions blow up in finite time; if $x_{0}=\sqrt{2/\pi}$, $x(t)\rightarrow\infty$ as $t\rightarrow\infty$, and finally if $x_{0}<\sqrt{2/\pi}$, $x(t)\rightarrow 0$ as $t\rightarrow\infty$. We call the solution (2.2) with $x_{0}=\sqrt{2/\pi}$ the critical solution and denote it by $x_{c}(t)$.
@@
Given the explicit form (2.2) of solutions of (2.1), the asymptotics of the critical solution as $t\rightarrow\infty$ is easily computed; it is given by
$x_{c}(t)\sim t+\frac{1}{t}-\frac{2}{t^{3}}+O\left(\frac{1}{t^{5}}\right).$ (2.3)
Note that in terms of rivers, $x_{c}(t)$ provides a repelling river, while $x(t)\equiv 0$ and all the positive semi-orbits that converge to it, constitute attracting rivers.
The asymptotic expansion of the repelling river in (2.3) can be constructed directly from the equation (2.1) without using (2.2) by following the procedure explained in [1].
@@ 3 The linear case
$\text{\rm{d}}X(t)=\big{(}cX(t)+f(t)\big{)}\,\text{\rm{d}}t+\sigma\text{\rm{d}}W(t),\;t\geq 0,$ (3.1)
@@
The SDE (3.1) with initial condition $X(0)=x$ has the explicit solution
@@
It follows that all trajectories are exponentially unstable with $\frac{\text{\rm{d}}}{\text{\rm{d}}x}~{}X(t,x)~{}=~{}\text{\rm{e}}^{ct}$, $t\geq 0,\,x\in\mathbb{R}$. We now assume that
@@
Below we will be using improper Itô integrals. For a rigorous definition of this mathematical object see, e.g., [3, Definition 25.10]. Denote the unique solution $X(t),\,t\geq 0$ of (3.1) with the initial condition
@@
by $\mathscr{R}(t),\,t\geq 0$ (here and below $\mathscr{R}$ stands for river). Then
$X(t,X(0))$ $:=\mathscr{R}(t)=-\text{\rm{e}}^{ct}\int_{t}^{\infty}\text{\rm{e}}^{-cs}f(s)\,\text{\rm{d}}s-\sigma e^{ct}\int_{t}^{\infty}\text{\rm{e}}^{-cs}\,\text{\rm{d}}W(s)$
$=-\text{\rm{e}}^{ct}\int_{t}^{\infty}\text{\rm{e}}^{-cs}f(s)\,\text{\rm{d}}s-\sigma A(t),$
@@
so solutions go to +$\infty$+ ∞ (respectively, $\infty$- ∞) exponentially fast with rate $c$ when $x$ is above (respectively below) $\mathscr{R}(0)$ while if $x=\mathscr{R}(0)$,
$\big{|}X(t,\mathscr{R}(0))\big{|}=\big{|}\mathscr{R}(t)\big{|}=o\big{(}\text{\rm{e}}^{ct}\big{)}.$
In particular, if in (3.1) we set $f\equiv 0$, then just like in the deterministic case ($\sigma\equiv 0$) where the repelling river given $X(t,x)\equiv 0$, we have a trichotomy with random repelling river given by $\mathscr{R}(t)$.
@@ 4 Main results
We now study the additive noise stochastic equivalent of (2.1), namely, the stochastic differential equation
$\text{\rm{d}}X(t)=X(t)\big{(}X(t)-t\big{)}\,\text{\rm{d}}t+\sigma\,\text{\rm{d}}W(t),\,X(s)=x,\,t\geq s,$ (4.1)
@@
Before stating our first main result, we collect some properties of the following more general model:
$X(t)=x+\int_{s}^{t}X(u)\big{(}X(u)-u\big{)}\,\text{\rm{d}}u+H(t)-H(s),\,t\geq s,$ (4.2)
where $H:[0,\infty)\to\mathbb{R}$ is a given (random or deterministic) continuous function satisfying $H(0)=0$, $x\in\mathbb{R}$, and $s\geq 0$. Again we denote the unique local solution of (4.2) by $X_{s,t}(x)$.
First of all, arguing from the contrary from (4.2), we have the fundamental monotonicity property:
Lemma 4.1.
$x<y$ implies $X_{s,t}(x)\leq X_{s,t}(y)$ for $t\geq s$; moreover, $X_{s,t}(x)<X_{s,t}(y)$ when $\beta_{s}(x)>t$.
We define the blow-up time by
@@
Proposition 4.2.
For each $x\in\mathbb{R}$ and $v\geq s\geq 0$, we have
$\inf_{s\leq t\leq v}X_{s,t}(x)\geq x\wedge 0+\inf_{s\leq t\leq u\leq v}\big{(}H(u)-H(t)\big{)}.$
Proof.
Assume that $x\leq 0$. Then
$\inf_{s\leq t\leq v}$ $X_{s,t}(x)\geq\inf_{s\leq t\leq v}\left(x+\int_{s}^{t}X(u)\big{(}X(u)-u\big{)}\,\text{\rm{d}}u\right)+\inf_{s\leq t\leq v}\left(H(t)-H(s)\right)$
$\geq$ $x\wedge 0+\inf_{s\leq t\leq u\leq v}\big{(}H(u)-H(t)\big{)},$
where the inequality
$\inf_{s\leq t\leq v}\left(x+\int_{s}^{t}X(u)\big{(}X(u)-u\big{)}\,\text{\rm{d}}u\right)\geq x\wedge 0$
follows since $y(y-r)\geq 0$ for $r\geq 0$ and $y\leq 0$. The case of $x>0$ follows by monotonicity of Lemma 4.1. ∎
Proposition 4.2 implies that the function $t\mapsto X_{s,t}(x)$ cannot reach $\infty$- ∞ in finite time. Further, if $X_{s,t}(x)\geq A$ for some $t\in[s,v]$ and $A\geq v+\inf_{s\leq\kappa\leq u\leq v}\big{(}H(u)-H(\kappa)\big{)}$, then
@@
by the same argument as was used in Proposition 4.2, since $y(y-r)\geq 0$ for $y\geq v$ and $r\in[s,v]$. This result implies that $\lim_{t\uparrow\beta_{s}(x)}X_{s,t}(x)=\infty$ whenever $\beta_{s}(x)<\infty$. The function $x\mapsto\beta_{s}(x)$ is non-decreasing for each $s\geq 0$.
@@
Proposition 4.3.
For any $s\geq 0$, we have $\beta_{s}(x)<\infty$ for sufficiently large $x$.
Proof.
To see this, let $H_{*}$ be a lower bound of $H(t)-H(s)$ for $t\in[s,s+1]$ and let $x$ be so large that the solution of $Y^{\prime}=\frac{1}{2}Y^{2}$ with initial condition $x+H_{*}-1$ at time $s$ blows up before time $s+1$. In addition, assume that $x\geq 3-H_{*}$. Let $X(t):=X_{s,t}(x)$, $t\geq s$. Then $X(s)>Y(s)$ since $H_{*}\leq 0$. Let $\tau$ be the infimum of all $t\geq s$ such that $X(t)=Y(t)$ and assume that $\tau\leq s+1$. For $u\leq\tau$ we have $X(u)\geq Y(u)\geq x+H_{*}-1$ since $Y$ is increasing. Then, for $u\in[s,\tau]$,
$X(u)$ $\big{(}X(u)-1\big{)}=\frac{1}{2}X^{2}(u)+\frac{1}{2}X(u)\big{(}X(u)-2\big{)}\geq\frac{1}{2}Y^{2}(u)$
$+\frac{1}{2}X(u)\big{(}X(u)-2\big{)}\geq\frac{1}{2}Y^{2}(u),$
since $X(u)-2\geq Y(u)-2\geq 0$. Hence
$X(\tau)$ $=x+\int_{s}^{\tau}X(u)\big{(}X(u)-1\big{)}\,\text{\rm{d}}u+H(\tau)-H(s)$
$>x+H_{*}-1+\frac{1}{2}\int_{s}^{\tau}Y^{2}(u)\text{\rm{d}}u=Y(\tau),$
which is a contradiction. ∎
From Proposition 4.3 it follows by monotonicity that the set
$I_{1}(s):=\{x\in\mathbb{R}\,|\,\beta_{s}(x)<\infty\}$ (4.3)
is an interval which is unbounded above and non-empty.
Now we define
$x_{\inf}(s):=\inf\{x\in\mathbb{R}:\,\beta_{s}(x)<\infty\}\in[-\infty,\infty).$
Proposition 4.4.
$x_{\inf}(t)=X_{s,t}(x_{\inf}(s))$
whenever $s>t$ and $x_{\inf}(s)>-\infty$.
Proof.
Assume that for some $t>s$ we have that $X_{s,t}(x_{\inf}(s))<x_{\inf}(t)$. Consider $x\in(X_{s,t}(x_{\inf}(s)),x_{\inf}(t))$. Then by monotonicity there exists $y$ such that $x=X_{s,t}(y)$ and $y>x_{\inf}(s)$. Hence $X_{s,t}(y)$ blows up in finite time but since $x<x_{\inf}(t)$, this is impossible. The argument for the case $X_{s,t}(x_{\inf}(s))>x_{\inf}(t)$ is similar. ∎
It is obvious from Proposition 4.4 that $X_{s,t}(x_{\inf}(s))$ is defined for all $t>s$.
In particular, the set $I_{1}(s)$ of (4.3) is open; depending on $H$ it may be all of $\mathbb{R}$ or a proper (non-empty) subset of $\mathbb{R}$.
@@
Theorem 4.5 implies that for each $s\geq 0$ and $x\in\mathbb{R}$ the set $\Omega_{s,x}\subset\Omega$ consisting of those $\omega\in\Omega$ for which the solution starting at $(s,x)$ neither blows up nor converges to zero has measure 0. For $s\geq 0$ denote the union of all $\Omega_{s,x},\,x\in\mathbb{R}$ by $\Omega_{s}$. Since this is an uncountable union it does not follow that $\Omega_{s}$ has measure 0. Indeed we will see in the following that $\mathbb{P}(\Omega_{s})>0$ for every $s\geq 0$. In order to obtain a better understanding of the set $\Omega_{s}$ we will use the notation $I_{1}(s,\omega)$ of those $x\in\mathbb{R}$ for which $\beta_{s}(x)<\infty$ corresponding to a realisation $\omega$, consistent with (4.3). We saw that $I_{1}(s,\omega)$ is a non-empty open interval which is unbounded from above. Similarly, we set
$I_{2}(s,\omega):=\{x\in\mathbb{R}\,|\,X_{s,t}(x)\rightarrow 0\hbox{ as }t\rightarrow\infty\}.$ (4.4)
Again, by monotonicity of solutions, $I_{2}(s,\omega)$ is a (possibly empty) interval. Theorem 4.5 implies that every $x\in\mathbb{R}$ is almost surely contained in the (disjoint) union of these two intervals, so the complement of the union of these intervals can contain at most one point. We set
$\mathscr{R}(s,\omega)=x_{\inf}(s)$
for a realisation $\omega\in\Omega$, i.e. it is the infimum of $I_{1}(s,\omega)$ (which necessarily coincides with the supremum of $I_{2}(s,\omega)$). Note that $\mathscr{R}(s,\omega)=-\infty$ iff $I_{2}(s,\omega)$ is empty iff $I_{1}(s,\omega)=\mathbb{R}$. We know that $\mathscr{R}(s,\omega)$ does not belong to $I_{1}(s,\omega)$ but at the moment, it is unclear if $\mathscr{R}(s,\omega)$ belongs to $I_{2}(s,\omega)$. We will see later that it almost surely doesn’t. Therefore (almost surely), $\omega\in\Omega_{s}$ iff $\mathscr{R}(s,\omega)>-\infty$ and $\omega\in\Omega_{s,x}$ iff $\mathscr{R}(s,\omega)=x$. In particular, $\mathbb{P}(\Omega_{s})=\mathbb{P}(\mathscr{R}(s,\omega)>-\infty)$ and we will see in Theorem 4.9 that this quantity converges to 1 as $s\to\infty$.
@@
Recall that $\mathscr{R}(s)\in[-\infty,\infty)$ for all $s\geq 0$ and that $\mathscr{R}(s)=-\infty$ implies $\mathscr{R}(r)=-\infty$ for every $r<s$. Further, $X_{s,t}(\mathscr{R}(s))=\mathscr{R}(t)$ whenever $t\geq s$ and $\mathscr{R}(s)>-\infty$, so $\mathscr{R}(t)$ is a solution of (4.1) on the interval where it is larger than $\infty$- ∞. Theorem 4.7 implies that, for each $\varepsilon>0$ and $\alpha\in(0,1/2)$, $\lim_{s\to\infty}\mathbb{P}\big{(}s^{\alpha}|\mathscr{R}(s)-s|\geq\varepsilon\big{)}=0$ but the following result is stronger.
@@
In the deterministic case, it follows from the fact that the initial value of the repelling river at time 0 is positive, together with the fact that the derivative $x^{\prime}$ in (2.1) is zero on the diagonal $x=t$ that the function $\mathscr{R}$ lies above the diagonal, i.e. $\mathscr{R}(s)>s$ for all $s\geq 0$. This is not true in the stochastic case.
@@
$\big{|}\mathscr{R}_{n}(t)-\mathscr{R}(t)\big{|}=o(t^{-\frac{3}{2}n-\frac{1}{2}+\varepsilon}),\;t\to\infty,$ (4.5)
@@
Define $\Gamma_{n}(t):=\sup_{u\geq t}\big{|}\mathscr{R}_{n}(u)-\mathscr{R}(u)\big{|}$. Choose $\varepsilon<1/2$. By Theorem 4.9, there exists some $t_{0}(\varepsilon,\omega)<\infty$ such that $|Z(u)|\leq u^{-\frac{1}{2}+\varepsilon}$ for all $u\geq t_{0}(\varepsilon,\omega)$. Hence, for $t\geq t_{0}(\varepsilon,\omega)$,
@@
Let $t\geq t_{1}(\varepsilon,\omega):=3\vee t_{0}(\varepsilon,\omega)$. Since $\varepsilon<1/2$, by induction, $\Gamma_{n}(t)\leq t^{-\frac{1}{2}+\varepsilon}$ (this is true for $n=1$) since
@@
  • The solution with initial condition $(s,x)$ with $x\in(0,s-1]$ reaches the set $\{0,s\}$ in finite time. Let $\rho(s,x)$ be the probability that it hits level 0 before level $s$. Lemma 5.7 says that $\lim_{s\to\infty}\inf_{x\leq s-1}\rho(s,x)=1$ which implies that $\overline{\rho}(s):=\inf_{t\geq s}\inf_{x\in(0,s-1]}\rho(t,x)$ converges to 1 as well.
  • The solution with initial condition $(s,x)$, $x<0$ will hit 00 after finite time (irrespective of the starting time $s\geq 0$). This is Lemma 5.6.
@@
These statements together imply the claims in Theorems 4.5 and 4.7 as follows:
@@
$\inf_{x\in\mathbb{R}}\Big{(}B(s,x)+C(s,x)\Big{)}\geq\overline{B}(s)\wedge\big{(}\overline{\rho}(s)\overline{\chi}(s)\big{)}$ (5.1)
@@
$\kappa(s,x)\leq\int_{\mathbb{R}}\kappa(t,y)\,\mathbb{P}\big{(}X_{s,t}(x)\in\text{\rm{d}}y\big{)}$ (5.2)
@@
We will now formulate and prove the lemmas mentioned above. Since the proof of Theorem 4.9 will require finer estimates of $B(s,x)$ than those of Theorems 4.5 and 4.7 we will formulate the lemmas accordingly. The first lemma provides a lower bound for the blow-up probability when a trajectory starts slightly above the diagonal. The stated result is in fact more precise than needed to prove the two theorems. It will be convenient to work with the process
@@
$\text{\rm{d}}Y(t)=\big{(}Y(t)(Y(t)+t)-1\big{)}\,\text{\rm{d}}t+\sigma\,\text{\rm{d}}W(t),\,t\geq s,\,Y(s)=x.$ (5.3)
@@
$B\big{(}s,s+s^{-\alpha}\big{)}\geq 1-\exp\Big{\{}-\frac{1}{\sigma^{2}}s^{1-2\alpha}\big{(}1+o(1)\big{)}\Big{\}},\,s\to\infty.$
@@
$\text{\rm{d}}\,\overline{Y}(t)=\big{(}\overline{Y}(t)(\overline{Y}(t)+s)-1\big{)}\,\text{\rm{d}}t+\sigma\,\text{\rm{d}}W(t),\,t\geq s,\overline{Y}(s)=\gamma.$ (5.4)
$\overline{Y}$ solves an equation with coefficients independent of $t$ and is therefore easier to analyze than $Y$. Note that $\overline{Y}(t)\leq Y(t)$ as long as $Y$ is non-negative. Further, $\overline{Y}$ blows up to $\infty$ with probability 1 by Feller’s test of explosion, see Lemma A.1c), applied to $(l,r)=\mathbb{R}$. Therefore,